# Degree two Gopakumar–Vafa invariants of local curves

### v3

@@ 1.1. Motivation and results
Let $X$ be a smooth projective CY $3$-fold. For a curve class $\beta\in H_{2}(X,\mathbb{Z})$ and a non-negative integer $g\in\mathbb{Z}_{\geq 0}$, let $\operatorname{GW}_{g,\beta}\in\mathbb{Q}$ (resp. $n_{g,\beta}\in\mathbb{Z}$) denote the GW (resp. GV) invariant of $X$. Then we have the equation:
@@
Let $C$ be a smooth projective curve of genus two, $L\in\operatorname{Pic}^{3}(C)$ a generic line bundle. Let $[N]\in\mathbb{P}\operatorname{Ext}^{1}(L,L^{-1}\otimes\omega_{C})$ be a generic rank two vector bundle on $C$, and let $X=\operatorname{Tot}_{C}(N)$. Then we have
@@
Let $C$ be a smooth projective curve of arbitrary genus, and let $L\in\operatorname{Pic}^{2g(C)-1}(C)$ be a generic line bundle. Let $[N]\in\mathbb{P}\operatorname{Ext}^{1}(L,L^{-1}\otimes\omega_{C})$ be a generic rank two vector bundle on $C$, and let $X=\operatorname{Tot}_{C}(N)$. Then we have
@@ 1.2. GV invariants of a local curve and strategy of the proofs
First, we review the proposed definition of Maulik and Toda. Let $X$ be a smooth quasi-projective CY $3$-fold. We fix a curve class $\beta\in H_{2}(X,\mathbb{Z})$. Let $M_{X}(\beta,1)$ be the moduli space of slope stable (with respect to a fixed ample divisor) one-dimensional sheaves $E$ on $X$ with $[E]=\beta$ and $\chi(E)=1$. We have the Hilbert–Chow morphism:
@@
When $r=2$, we will give a general strategy to compute both (1) and (2). For (1), we will use the explicit description of $\operatorname{H}^{*}(M_{L}(2,1),\mathbb{Q})$ by generators and relations [dCHM12, HT03, HT04].
For (2), we will calculate the fibrewise contributions and then integrate. Namely, we have a stratification on $U_{\varepsilon}$ given by the types of singularities of the spectral curves. Since the nearby hyperplane $U_{\varepsilon}$ avoids the origin and we assume $r=2$, the spectral curves $\widetilde{C}$ over points in $U_{\varepsilon}$ are all reduced (more precisely, we first need to remove the trace part $\operatorname{H}^{0}(L)$). Hence, by [MS13, MY14], the fibrewise contribution can be computed by the Euler characteristics of the Hilbert schemes $\operatorname{Hilb}^{n}(\widetilde{C})$. The latter is then computed by certain HOMFLY polynomials [Mau16, OS12]. It remains to compute the Euler characteristic of each stratum $S\subset U_{\varepsilon}$. Our main idea is to realise (a resolution of) the closure of $S$ as a degeneracy locus inside a product of $C$. These varieties are closely related to generalized de Jonquières divisors [Far23, Ung21].
@@ 1.3. Relation to previous work
  • $X=\operatorname{Tot}_{C}(\omega_{C}\oplus\mathcal{O}_{C})$ and $\beta$ is arbitrary [CDP14, HMMS22, MS24],
@@
For $X=\operatorname{Tot}_{C}(\omega_{C}\oplus\mathcal{O}_{C})$, as observed in [CDP14], the GV/PT correspondence follows from the P=W conjecture, which has recently been proved [HMMS22, MS24].
@@ 2.1. Twisted Higgs bundles
for some line bundle $L$ with $\deg(L)\geq 2g(C)-1$. Let $f\colon\operatorname{H}^{0}(L^{\otimes 2})\to\mathbb{C}$ be a linear function corresponding to the class $[N]\in\operatorname{Ext}^{1}(L,L^{-1}\otimes\omega_{C})$ via Serre duality: $Ext^{1}(L,L^{-1}\otimes\omega_{C})\cong\operatorname{H}^{0}(L^{\otimes 2})^{\vee}$. Then we have an isomorphism
@@ 2.2. Rigidity of vector bundles
Fix a line bundle $L$ on $C$ with degree $d\geq 2g(C)-1$. Assume that $L$ is a generic element in $\operatorname{Pic}^{d}(C)$. Then a general $[N]\in\mathbb{P}\operatorname{Ext}^{1}(L,L^{-1}\otimes\omega_{C})$ is $2$-rigid.
@@
for $10\leq d\leq g(C)-$ 1.
It remains to show that a generic $[N]\in\mathbb{P}\operatorname{Ext}^{1}(L,L^{-1}\otimes\omega_{C})$ is slope semistable, which follows from Lemma 2.7 below. ∎
An earlier version of this paper omitted the check that generic $[N]\in\mathbb{P}\operatorname{Ext}^{1}(L,L^{-1}\otimes\omega_{C})$ is slope semistable; we thank the anonymous referee for pointing this out, as well as suggesting the proof of the following lemma.
@@
Let $L$ be a line bundle of degree $d\geq g(C)-1$. Assume that $L$ is a generic element in $\operatorname{Pic}^{d}(C)$. Then a general $[N]\in\mathbb{P}\operatorname{Ext}^{1}(L,L^{-1}\otimes\omega_{C})$ is $\mu$-semistable.
@@
Since $\mu$-semistability is an open condition, it is enough to find a line bundle $L\in\operatorname{Pic}^{d}(C)$ and a point $[N]\in\mathbb{P}\operatorname{Ext}^{1}(L,L^{-1}\otimes\omega_{C})$ such that $N$ is $\mu$-semistable.
We may find a line bundle $L_{0}\in\operatorname{Pic}^{d}(C)$ and line bundles $M_{0},M_{1}$ of degree $g(C)-1$ such that there is a surjection $M_{0}\oplus M_{1}\to L_{0}$. We may then take a line bundle $L^{\prime}\in\operatorname{Pic}^{0}(C)$ such that $M_{0}\otimes M_{1}\otimes(L^{\prime})^{\otimes 2}\cong\omega_{C}$. Now, set $N\coloneqq(M_{0}\oplus M_{1})\otimes L^{\prime}$ and $L\coloneqq L_{0}\otimes L^{\prime}$. Since $M_{0}$ and $M_{1}$ have the same degree, $N$ is $\mu$-semistable. Furthermore, we have $L\in\operatorname{Pic}^{d}(C)$ and there exists a short exact sequence
@@ 2.3. Hitchin fibrations and Gopakumar–Vafa invariants
In this subsection, we fix a line bundle $L$ on $C$ with $\deg(L)\geq 2g(C)-1$ and an extension class $[N]\in\mathbb{P}\operatorname{Ext}^{1}(L,L^{-1}\otimes\omega_{C})$. By the Serre duality $\operatorname{Ext}^{1}(L,L^{-1}\otimes\omega_{C})\cong\operatorname{H}^{0}(L^{\otimes 2})^{\vee}$, the class $[N]$ defines a linear function $f\colon\operatorname{H}^{0}(L^{\otimes 2})\to\mathbb{C}$. To simplify the notation, we put $M_{L}(r,1)\coloneqq M^{\mathrm{st}}_{L}(r,1)$. We will also consider the moduli space of traceless Higgs bundles:
@@
[…]
@@
$M_{L}(r,1)\xrightarrow{\sim}\operatorname{H}^{0}(L)\times M^{0}_{L}(r,1),\quad(E,\phi)\mapsto(\operatorname{tr}(\phi),(E,\phi-\operatorname{tr}(\phi)\operatorname{id})).$
@@
where $P_{0}$ denotes the stalk at the origin, and $U_{\varepsilon}\coloneqq f^{-1}(\varepsilon)$ ($0<\varepsilon\ll 1$) is the nearby hyperplane.
@@
$\operatorname{H}(B_{L},\phi_{f}(P))\rightarrow\operatorname{H}(B_{L},P\lvert_{f^{-1}(0)})\rightarrow\operatorname{H}(B_{L},\psi_{f}P)\rightarrow$
it is sufficient to prove that $\chi(\psi_{f}P)=\chi(P\lvert_{U_{\epsilon}})$. By Lemma 2.11 again, this is equivalent to showing that
$\chi((\psi_{f}P)_{0})=\chi(P\lvert_{U_{\epsilon}})$
@@
For $[N]\in\mathbb{P}\operatorname{Ext}^{1}(L,L^{-1}\otimes\omega_{C})$, we have
@@
    Computation of $n_{g,2[C]}(\operatorname{Tot}_{C}(L))$: this can be done by using the explicit description of the global cohomology $\operatorname{H}^{*}(M_{L}(2,1),\mathbb{Q})$ by generators and relations [HT03, HT04]. The cohomological and the perverse degree of each generator is determined in [dCHM12].
@@ 3. Genus two
Let $C$ be a smooth projective curve of genus two. Let $L$ be a line bundle on $C$ of degree $13=2g(C)-$ 1. Let $f\colon\operatorname{H}^{0}(L^{\otimes 2})\to\mathbb{C}$ be a linear function. The goal of this section is to compute the GV invariants associated to $L$ and $f$. We apply the strategy explained in Remark 2.16.
@@ 3.1. Stratification of the Hitchin base
    $d_{\lambda}\cdot e(S_{\lambda}\cap U_{\varepsilon})=e(Z)-e(D)-\sum_{\mu\vdash\lambda}d_{\mu,\lambda}\cdot e(S_{\mu}\cap U_{\varepsilon})$
@@
    Suppose that $\mathbb{P}\operatorname{H}^{0}(L^{\otimes 2})\cap S_{\lambda}$ is 00-dimensional. Since $\mathbb{P}\operatorname{H}^{0}(L^{\otimes 2})\subset\operatorname{Sym}^{2g(C)-1}(C)$ has codimension $g=\dim J(C)$ and the dimension of $S_{\lambda}$ is equal to the number of rows in $\lambda$, this happens precisely when
    #{rows in $\#\{\text{rows in }\lambda\}=g(C).$
@@
    $e(U_{\varepsilon}\cap S_{\lambda})=e(Z)/d_{\lambda},$
    is given by the classical de Jonquières formula [dJ66].
@@
    Since $k>g(C)$, the Abel–Jacobi map $\operatorname{Sym}^{k}(C)\to\operatorname{Pic}^{k}(C)$ is a projective space bundle, hence so is $Z^{\prime}\to C^{\times(m-k)}$. It is easier to compute $e(Z^{\prime})$ than to compute $e(Z)$ for a degeneracy locus $Z$.
@@
(3.2) $e(U_{\varepsilon}\cap S_{\lambda})=\begin{cases}0&(\lambda=(6)),\\ 50&(\lambda=(5,1)),\\ 128&(\lambda=(4,2)),\\ -216&(\lambda=(4,1,1)),\\ 81&(\lambda=(3,3)),\\ -668&(\lambda=(3,2,1)),\\ 542&(\lambda=(3,1,1,1)),\\ -128&(\lambda=(2,2,2)),\\ 968&(\lambda=(2,2,1,1)),\\ -1012&(\lambda=(2,1,1,1,1)).\end{cases}$
@@ 3.1.2. $\lambda=(5,1)$
[…]
@@ 3.1.4. $\lambda=(4,1,1)$
as follows. The divisor $H_{\infty}$ is linearly equivalent to the divisor $D_{p}$ defined by the vanishing of the section $s$ at $p$, for some fixed $p\in C$. If a section $s=\psi(q)$ in $Z$ vanishes at $p$, either $p=x$, in which case the intersection multiplicity is $4$, since the regular function defining $D_{p}$ vanishes to order 4,or $p$ is one of the other two zeros of $\Psi(q)$. Counting these points corresponds to counting solutions to $L^{\otimes 2}(-p)\cong\mathcal{O}_{C}(4x+y)$. But now this is the same theta divisor calculation as we have seen before, and we find that there are $24^{2}*1^{2}*$ 2 such points. So $Z\cap H_{\infty}=32+4=36$.
@@ 3.1.5. $\lambda=(3,3)$
$1623^{2}\cdot 3^{2}\cdot\Theta^{2}=$ 162 .
@@ 3.1.6. $\lambda=(3,2,1)$
Let $Z\subset C\times C\times C$ be the subvariety defined by $\mathcal{O}_{C}(3x+2y+z)\cong L^{\otimes 2}$. We claim that $e(Z)=-196$. We postpone the proof until the next subsection, see Proposition 3.6.
Fixing some $p\in C$, the preimage $D$ of the divisor $H_{\infty}$ is equivalent to the divisor $3\{p\}\times C\times C+2C\times\{p\}\times C+C\times C\times\{p\}$. So $D$ consists of $132=3*8+2*18+72$ points. These numbers come from the same calculation of intersections of theta divisors as before, so for example $72$ appears via $72=2*2^{2}*3^{2}$. $Z$ also contains some copies of the deeper strata: it contains one copy of $S_{(4,2)}$ (corresponding to $z=x$), one copy of $S_{(5,1)}$ given by setting $x=y$, and a variety $S^{[2\operatorname{\!\colon\!\!\!}1]}_{(3,3)}$ (corresponding to $y=z$). So putting everything together we find
$e(S_{(3,2,1)}\cap U_{\varepsilon})=-196-132-128-50-162=-668.$

### v4

@@ 1.1. Motivation and results
Let $X$ be a smooth projective CY $3$-fold. For a curve class $\beta\in H_{2}(X,\mathbb{Z})$ and a non-negative integer $g\in\mathbb{Z}_{\geq 0}$, let $\operatorname{GW}_{g,\beta}\in\mathbb{Q}$ (resp. $n_{g,\beta}\in\mathbb{Z}$) denote the GW (resp. GV) invariant of $X$. Then
@@
Let $C$ be a smooth projective curve of genus two, $L\in\operatorname{Pic}^{3}(C)$ a generic line bundle. Let $N]\in\mathbb{P}\operatorname{Ext}^{1}(L,L^{-1}\otimes\omega_{C})$ be a generic rank two vector bundle on $C$, and let $X=\operatorname{Tot}_{C}(N)$. Then we have
@@
Let $C$ be a smooth projective curve of arbitrary genus, and let $L\in\operatorname{Pic}^{2g(C)-1}(C)$ be a generic line bundle. Let $N]\in\mathbb{P}\operatorname{Ext}^{1}(L,L^{-1}\otimes\omega_{C})$ be a generic rank two vector bundle on $C$, and let $X=\operatorname{Tot}_{C}(N)$. Then
@@ 1.2. GV invariants of a local curve and strategy of the proofs
First, we review the proposed definition of Maulik and Toda. Let $X$ be a smooth quasi-projective CY $3$-fold. We fix a curve class $\beta\in H_{2}(X,\mathbb{Z})$. Let $M_{X}(\beta,1)$ be the moduli space of slope stable (with respect to a fixed ample divisor) one-dimensional sheaves $E$ on $X$ with $E]=\beta$ and $\chi(E)=1$. We have the Hilbert–Chow morphism:
@@
When $r=2$, we will give a general strategy to compute both (1) and (2). For (1), we will use the explicit description of $\operatorname{H}^{*}(M_{L}(2,1),\mathbb{Q})$ by generators and relations [dHM12, HT03, HT04].
[…]
@@ 1.3. Relation to previous work
  • $X=\operatorname{Tot}_{C}(\omega_{C}\oplus\mathcal{O}_{C})$ and $\beta$ is arbitrary [CDP14, HMM+22, MS24],
@@
For $X=\operatorname{Tot}_{C}(\omega_{C}\oplus\mathcal{O}_{C})$, as observed in [CDP14], the GV/PT correspondence follows from the P=W conjecture, which has recently been proved [HMM+22, MS24].
@@ 2.1. Twisted Higgs bundles
for some line bundle $L$ with $\deg(L)\geq 2g(C)-1$. Let $f\colon\operatorname{H}^{0}(L^{\otimes 2})\to\mathbb{C}$ be a linear function corresponding to the class $N]\in\operatorname{Ext}^{1}(L,L^{-1}\otimes\omega_{C})$ via Serre duality: $Ext^{1}(L,L^{-1}\otimes\omega_{C})\cong\operatorname{H}^{0}(L^{\otimes 2})^{\vee}$. Then we have an isomorphism
@@ 2.2. Rigidity of vector bundles
Fix a line bundle $L$ on $C$ with degree $d\geq 2g(C)-1$. Assume that $L$ is a generic element in $\operatorname{Pic}^{d}(C)$. Then a general $N]\in\mathbb{P}\operatorname{Ext}^{1}(L,L^{-1}\otimes\omega_{C})$ is $2$-rigid.
@@
for $0\leq d\leq g(C)-1$.
It remains to show that a generic $N]\in\mathbb{P}\operatorname{Ext}^{1}(L,L^{-1}\otimes\omega_{C})$ is slope semistable, which follows from Lemma 2.7 below. ∎
An earlier version of this paper omitted the check that generic $N]\in\mathbb{P}\operatorname{Ext}^{1}(L,L^{-1}\otimes\omega_{C})$ is slope semistable; we thank the anonymous referee for pointing this out, as well as suggesting the proof of the following lemma.
@@
Let $L$ be a line bundle of degree $d\geq g(C)-1$. Assume that $L$ is a generic element in $\operatorname{Pic}^{d}(C)$. Then a general $N]\in\mathbb{P}\operatorname{Ext}^{1}(L,L^{-1}\otimes\omega_{C})$ is $\mu$-semistable.
@@
Since $\mu$-semistability is an open condition, it is enough to find a line bundle $L\in\operatorname{Pic}^{d}(C)$ and a point $N]\in\mathbb{P}\operatorname{Ext}^{1}(L,L^{-1}\otimes\omega_{C})$ such that $N$ is $\mu$-semistable.
We may find a line bundle $L_{0}\in\operatorname{Pic}^{d}(C)$ and line bundles $M_{0},M_{1}$ of degree g(C)1g(C)-1 such that there is a surjection $M_{0}\oplus M_{1}\to L_{0}$. We may then take a line bundle $L^{\prime}\in\operatorname{Pic}^{0}(C)$ such that $M_{0}\otimes M_{1}\otimes(L^{\prime})^{\otimes 2}\cong\omega_{C}$. Now, set $N\coloneqq(M_{0}\oplus M_{1})\otimes L^{\prime}$ and $L\coloneqq L_{0}\otimes L^{\prime}$. Since $M_{0}$ and $M_{1}$ have the same degree, $N$ is $\mu$-semistable. Furthermore, we have $L\in\operatorname{Pic}^{d}(C)$ and there exists a short exact sequence
@@ 2.3. Hitchin fibrations and Gopakumar–Vafa invariants
In this subsection, we fix a line bundle $L$ on $C$ with $\deg(L)\geq 2g(C)-1$ and an extension class $N]\in\mathbb{P}\operatorname{Ext}^{1}(L,L^{-1}\otimes\omega_{C})$. By the Serre duality $\operatorname{Ext}^{1}(L,L^{-1}\otimes\omega_{C})\cong\operatorname{H}^{0}(L^{\otimes 2})^{\vee}$, the class $[N]$ defines a linear function $f\colon\operatorname{H}^{0}(L^{\otimes 2})\to\mathbb{C}$. To simplify the notation, we put $M_{L}(r,1)\coloneqq M^{\mathrm{st}}_{L}(r,1)$. We will also consider the moduli space of traceless Higgs bundles:
@@
$\tilde{f}\colon\widetilde{B}_{L}\to\operatorname{H}^{0}(L^{\otimes 2})\xrightarrow{f}\mathbb{C},$
@@
$M_{L}(r,1)\xrightarrow{\sim}\operatorname{H}^{0}(L)\times M^{0}_{L}(r,1),\quad(E,\phi)\mapsto(\operatorname{tr}(\phi),(E,\phi-(\operatorname{tr}(\phi)/2)\operatorname{id})).$
@@
where $P_{0}$ denotes the stalk at the origin, and $U_{\varepsilon}\coloneqq f^{-1}(\varepsilon)$ (for $0<\varepsilon\ll 1$) is the nearby hyperplane.
@@
$\operatorname{H}(B_{L},\phi_{f}(P))\rightarrow\operatorname{H}(B_{L},P\lvert_{f^{-1}(0)})\rightarrow\operatorname{H}(B_{L},\psi_{f}(P))\rightarrow$
it is sufficient to prove that $\chi(\psi_{f}(P))=\chi(P\lvert_{U_{\epsilon}})$. By Lemma 2.11 again, this is equivalent to showing that
$\chi((\psi_{f}(P))_{0})=\chi(P\lvert_{U_{\epsilon}})$
@@
For $N]\in\mathbb{P}\operatorname{Ext}^{1}(L,L^{-1}\otimes\omega_{C})$, we have
@@
    Computation of $n_{g,2[C]}(\operatorname{Tot}_{C}(L))$: this can be done by using the explicit description of the global cohomology $\operatorname{H}^{*}(M_{L}(2,1),\mathbb{Q})$ by generators and relations [HT03, HT04]. The cohomological and the perverse degree of each generator is determined in [dHM12].
@@ 3. Genus two
Let $C$ be a smooth projective curve of genus two. Let $L$ be a line bundle on $C$ of degree 3=2g(C)13=2g(C)-1. Let $f\colon\operatorname{H}^{0}(L^{\otimes 2})\to\mathbb{C}$ be a linear function. The goal of this section is to compute the GV invariants associated to $L$ and $f$. We apply the strategy explained in Remark 2.16.
@@ 3.1. Stratification of the Hitchin base
    $d_{\lambda}\cdot\chi(S_{\lambda}\cap U_{\varepsilon})=\chi(Z)-\chi(D)-\sum_{\mu\vdash\lambda}d_{\mu,\lambda}\cdot\chi(S_{\mu}\cap U_{\varepsilon})$
@@
    Suppose that $\mathbb{P}\operatorname{H}^{0}(L^{\otimes 2})\cap S_{\lambda}$ is 0-dimensional. Since $\mathbb{P}\operatorname{H}^{0}(L^{\otimes 2})\subset\operatorname{Sym}^{2g(C)-1}(C)$ has codimension $g=\dim J(C)$ and the dimension of $S_{\lambda}$ is equal to the number of rows in $\lambda$, this happens precisely when
    $\{\text{rows in }\lambda\}=g(C).$
@@
    $\chi(U_{\varepsilon}\cap S_{\lambda})=\chi(Z)/d_{\lambda},$
    is given by the classical de Jonquières formula [DE 66].
@@
    Since $k>g(C)$, the Abel–Jacobi map $\operatorname{Sym}^{k}(C)\to\operatorname{Pic}^{k}(C)$ is a projective space bundle, hence so is $Z^{\prime}\to C^{\times(m-k)}$. It is easier to compute $\chi(Z^{\prime})$ than to compute $\chi(Z)$ for a degeneracy locus $Z$.
@@
(3.2) $\chi(U_{\varepsilon}\cap S_{\lambda})=\begin{cases}0&(\lambda=(6)),\\ 50&(\lambda=(5,1)),\\ 128&(\lambda=(4,2)),\\ -216&(\lambda=(4,1,1)),\\ 81&(\lambda=(3,3)),\\ -668&(\lambda=(3,2,1)),\\ 542&(\lambda=(3,1,1,1)),\\ -128&(\lambda=(2,2,2)),\\ 968&(\lambda=(2,2,1,1)),\\ -1012&(\lambda=(2,1,1,1,1)).\end{cases}$
@@ 3.1.2. $\lambda=(5,1)$
$\eta\colon\operatorname{Pic}^{1}(C)$$\operatorname{Pic}^{5}(C)$$\operatorname{Pic}^{1}(C).$$\scriptstyle{(-)^{\otimes 5}}$$\scriptstyle{L^{\otimes 2}\otimes(-)^{-1}}$
@@ 3.1.4. $\lambda=(4,1,1)$
as follows. The divisor $H_{\infty}$ is linearly equivalent to the divisor $D_{p}$ defined by the vanishing of the section $s$ at $p$, for some fixed $p\in C$. If a section $s=\psi(q)$ in $Z$ vanishes at $p$, either $p=x$, in which case the intersection multiplicity is $4$, since the regular function defining $D_{p}$ vanishes to order 4,or $p$ is one of the other two zeros of $\Psi(q)$. Counting these points corresponds to counting solutions to $L^{\otimes 2}(-p)\cong\mathcal{O}_{C}(4x+y)$. But now this is the same theta divisor calculation as we have seen before, and we find that there are $4^{2}*1^{2}*2$ such points. So $Z\cap H_{\infty}=32+4=36$.
@@ 3.1.5. $\lambda=(3,3)$
$3^{2}\cdot 3^{2}\cdot\Theta^{2}=162.$
@@ 3.1.6. $\lambda=(3,2,1)$
Let $Z\subset C\times C\times C$ be the subvariety defined by $\mathcal{O}_{C}(3x+2y+z)\cong L^{\otimes 2}$. We claim that $\chi(Z)=-196$. We postpone the proof until the next subsection, see Proposition 3.6.
Fixing some $p\in C$, the preimage $D$ of the divisor $H_{\infty}$ is equivalent to the divisor $3\{p\}\times C\times C+2C\times\{p\}\times C+C\times C\times\{p\}$. So $D$ consists of 132=38+218+72132=3*8+2*18+72 points. These numbers come from the same calculation of intersections of theta divisors as before, so for example $72$ appears via $72=2*2^{2}*3^{2}$. $Z$ also contains some copies of the deeper strata: it contains one copy of $S_{(4,2)}$ (corresponding to $z=x$), one copy of $S_{(5,1)}$ given by setting $x=y$, and a variety $S^{[2\operatorname{\!\colon\!\!\!}1]}_{(3,3)}$ (corresponding to $y=z$). So putting everything together we find
$\chi(S_{(3,2,1)}\cap U_{\varepsilon})=-196-132-128-50-162=-668.$